# Neural Network Diffusion

### v1

@@ A.1 Training recipe
A.2 Datasets
We evaluate the effectiveness of p-diff on 8 datasets. To be specific, CIFAR-10/100 (Krizhevsky et al., 2009). The CIFAR datasets comprise colored natural images of dimensions $32\times 32$, categorized into 10 and 100 classes, respectively. Each dataset consists of 50,000 images for training and 10,000 images for testing. ImageNet-1K (Deng et al., 2009) derived from the larger ImageNet-21K dataset, ImageNet-1K is a curated subset featuring 1,000 categories. It encompasses 1,281,167 training images and 50,000 validation images. STL-10 (Coates et al., 2011) comprises $96\times 96$ color images, spanning 10 different object categories. It serves as a versatile resource for various computer vision tasks, including image classification and object recognition. Flowers (Nilsback & Zisserman, 2008) is a dataset comprising 102 distinct flower categories, with each category representing a commonly occurring flower species found in the United Kingdom. Pets (Parkhi et al., 2012) includes around 7000 images with 37 categories. The images have large variations in scale, pose, and lighting. F-101 (Bossard et al., 2014) consists of 365K images that are crawled from Google, Bing, Yelp, and TripAdvisor using the Food-101 taxonomy.
In the appendix, we extend our p-diff in object detection, semantic segmentation, and image generation tasks. Therefore, we also introduce the extra-used datasets in the following. COCO (Lin et al., 2014) consists of over 200,000 images featuring complex scenes with 80 object categories. It is widely used for object detection and segmentation tasks. We implement image generation task on CIFAR-10.
A.3 Instructions for code
We have submitted the source code as the supplementary materials in a zipped file named as ‘p-diff.zip’ for reproduction. A README is also included for the instructions for running the code.
Appendix B Explorations of Designs and Strategies
@@ B.1 Why 1D CNNs?
B.2 Is variational autoencoder an alternative to our approach?
Variational autoencoder (VAE) (Cristianini et al., 2000) can be regarded as a probabilistic generative model and achieve many remarkable results in the generation area. We also implement VAE to generate neural network parameters. We first introduce the details of VAE in our experiment. We implement vanilla VAE using the same backbone of the autoencoder in p-diff for a fair comparison. We evaluate the VAE generator in the case of different $K$ and compare its best, average, and medium performances with p-diff generated models. Based on the results in Tab. 6(e), our approach outperforms VAE by a large margin in all cases. Another interesting finding is that the average performance of VAE generated models goes down as the number of original models increases.
B.3 Which normalization strategy is suitable?
@@
Appendix C More Ablations
@@
C.1 The intensity of noise added into input parameters
@@
C.2 The intensity of noise added into latent representations
Similar to Sec. C.1, we also ablate the noise intensity added into latent representations. As shown in Tab. 6(g), the performance stability of generated models becomes better as the noise intensity increases. However, too-large noise also breaks the distribution of the original latent representations.
C.3 The generalization on other types of parameters
@@
Appendix D Open Explorations
D.1 Do we need to train 1000-step diffusion model?
@@
D.2 Potential applications
Neural network diffusion can be utilized or help the following potential research areas. 1) Parameters initialization: our approach can generate high-performing initialized parameters. Therefore, that would speed up the optimization and reduce the overall cost of training. 2) Domain adaptation: our approach may have three benefits in the domain adaptation area. First, we can directly use the diffusion process to learn the well-performed models trained by different domain data. Second, some hard adaptations can be achieved by our approach. Third, the adaptation efficiency might be improved largely.
Appendix E Other Finding and Comparison Results
E.1 How to select generated parameters?
P-diff can rapidly generate numerous high-performance models. How do we evaluate these models? There are two primary strategies. The first one is to directly test them on the validation set and select the best-performing model. The second one is to compute the loss of model outputs compared to the ground truth on the training set to choose a model. We generated hundred model parameters with performance distributions in different intervals and displayed their accuracy curves on both the training and validation sets in Fig. 6(l). The experimental results indicate that p-diff exhibits a high level of consistency between the training and validation sets. To provide a fair comparison with baseline methods, we default to choose the model that performs the best results on the training set and compare it with the baseline.
(l) Accuracy distribution in p-diff models.
(m) Visualization of initial, SGD-trained, p-diff generated model.
@@ E.1 How to select generated parameters?
E.2 Parameter visualization
To provide a more intuitive understanding, we compare the parameters generated by our approach, SGD optimization (original), and randomly initialized. Taking ResNet-18 as an example, we report the mean, std, accuracy (acc.), and IoU of the normalization layer parameters of training on CIFAR-100 in Fig. 6(m). There is a significant difference between the parameters generated by our approach and the randomly initialized parameters, mean: 0.37 vs 0.36, std: 0.22 vs 0.21 The IoU between ours and SGD is 0.87. This visualization and results confirm that the diffusion process can learn the patterns of high-performance parameters and generate new good models from random noise. More importantly, our generated model has a great behavior contrast compared to the original model, which is reflected in the low IoU value.
E.3 Efficiency of parameter generation
@@
Appendix F Generalization on Other Tasks
@@
F.1 Object detection
Faster R-CNN (Ren et al., 2015) utilizes a region proposal network (RPN) which shares full-image convolutional features with the detection network to improve Fast R-CNN on object detection task. The FCOS (Fully Convolutional One-Stage) (Tian et al., 2020) model is a single-stage object detection model that simplifies the detection process by eliminating the need for anchor boxes. In the object detection task, We implement Faster R-CNN (Ren et al., 2015) and FCOS (Tian et al., 2020) with ResNet-50 backbone on the COCO (Lin et al., 2014) dataset based on torch/torchvision. Considering the time cost of data for p-diff, we directly use the pre-trained parameters as our first training data, then fine-tune it to obtain other training data. The parameters of the boxing predictor layer are generated by p-diff. We report the results in Tab. 10. Our method can get models with similar or even better performance than the original model in seconds.
@@ F.1 Object detection
F.2 Semantic segmentation
Fully Convolutional Network (FCN) (Long et al., 2015) was designed to efficiently process and analyze images at the pixel level, allowing for the semantic segmentation of objects within an image. Following the approach in object detection, we implement semantic segmentation task using FCN (Long et al., 2015) with ResNet-50 backbone to evaluate a subset of COCO val2017, on the 20 categories that are present in the Pascal VOC dataset. We generate a subset of the parameters of backbone and report the results in Tab. 11. Our approach can generate high-performing neural network parameters in semantic segmentation task.
@@ F.3 Image generation
DDPM (Ho et al., 2020) is a diffusion-based method in image generation, where UNet (Ronneberger et al., 2015) is used to model the noise. In the image generation task, we use p-diff to generate a subset of model parameters of UNet. For comparison, we evaluate the p-diff model’s FID score on the CIFAR-10 dataset and report the results in Tab. 12. The best p-diff generated UNet get similar performance to the original model.

### v2

@@ B.1 Training recipe
B.2 Datasets
We evaluate the effectiveness of p-diff on 8 datasets. To be specific, CIFAR-10/100 [36]. The CIFAR datasets comprise colored natural images of dimensions $32\times 32$, categorized into 10 and 100 classes, respectively. Each dataset consists of 50,000 images for training and 10,000 images for testing. ImageNet-1K [12] derived from the larger ImageNet-21K dataset, ImageNet-1K is a curated subset featuring 1,000 categories. It encompasses 1,281,167 training images and 50,000 validation images. STL-10 [10] comprises $96\times 96$ color images, spanning 10 different object categories. It serves as a versatile resource for various computer vision tasks, including image classification and object recognition. Flowers [50] is a dataset comprising 102 distinct flower categories, with each category representing a commonly occurring flower species found in the United Kingdom. Pets [51] includes around 7000 images with 37 categories. The images have large variations in scale, pose, and lighting. F-101 [4] consists of 365K images that are crawled from Google, Bing, Yelp, and TripAdvisor using the Food-101 taxonomy.
In the appendix, we extend our p-diff in object detection, semantic segmentation, and image generation tasks. Therefore, we also introduce the extra-used datasets in the following. COCO [41] consists of over 200,000 images featuring complex scenes with 80 object categories. It is widely used for object detection and segmentation tasks. We implement image generation task on CIFAR-10.
Appendix C Explorations of Designs and Strategies
@@ C.1 Why 1D CNNs?
C.2 Is variational autoencoder an alternative to our approach?
Variational autoencoder (VAE) [11] can be regarded as a probabilistic generative model and achieve many remarkable results in the generation area. We also implement VAE to generate neural network parameters. We first introduce the details of VAE in our experiment. We implement vanilla VAE using the same backbone of the autoencoder in p-diff for a fair comparison. We evaluate the VAE generator in the case of different $K$ and compare its best, average, and medium performances with p-diff generated models. Based on the results in Tab. 5(e), our approach outperforms VAE by a large margin in all cases. Another interesting finding is that the average performance of VAE generated models goes down as the number of original models increases.
C.3 Which normalization strategy is suitable?
@@
Appendix D More Ablations
@@
D.1 The intensity of noise added into input parameters
@@
D.2 The intensity of noise added into latent representations
Similar to Sec. D.1, we also ablate the noise intensity added into latent representations. As shown in Tab. 5(g), the performance stability of generated models becomes better as the noise intensity increases. However, too-large noise also breaks the distribution of the original latent representations.
D.3 The generalization on other types of parameters
@@
Appendix E Open Explorations
E.1 Do we need to train 1000-step diffusion model?
@@
E.2 Potential applications
Neural network parameter diffusion can be utilized or help the following potential research areas. 1) Parameters initialization: our approach can generate high-performing initialized parameters. Therefore, that would speed up the optimization and reduce the overall cost of training. 2) Domain adaptation: our approach may have three benefits in the domain adaptation area. First, we can directly use the diffusion process to learn the well-performed models trained by different domain data. Second, some hard adaptations can be achieved by our approach. Third, the adaptation efficiency might be improved largely.
Appendix F Other Finding and Comparison Results
F.1 Parameter patterns of original models.
F.2 How to select generated parameters?
P-diff can rapidly generate numerous high-performance models. How do we evaluate these models? There are two primary strategies. The first one is to directly test them on the validation set and select the best-performing model. The second one is to compute the loss of model outputs compared to the ground truth on the training set to choose a model. We generated hundred model parameters with performance distributions in different intervals and displayed their accuracy curves on both the training and validation sets in Fig. 6(a). The experimental results indicate that p-diff exhibits a high level of consistency between the training and validation sets. To provide a fair comparison with baseline methods, we default to choose the model that performs the best results on the training set and compare it with the baseline.
(a) Accuracy distribution in p-diff models.
(b) Visualization of initial, SGD-trained, p-diff generated model.
@@ F.2 How to select generated parameters?
F.3 Parameter visualization
To provide a more intuitive understanding, we compare the parameters generated by our approach, SGD optimization (original), and randomly initialized. Taking ResNet-18 as an example, we report the mean, std, accuracy (acc.), and IoU of the normalization layer parameters of training on CIFAR-100 in Fig. 6(b). There is a significant difference between the parameters generated by our approach and the randomly initialized parameters, mean: 0.37 vs 0.36, std: 0.22 vs 0.21 The IoU between ours and SGD is 0.87. This visualization and results confirm that the diffusion process can learn the patterns of high-performance parameters and generate new good models from random noise. More importantly, our generated model has a great behavior contrast compared to the original model, which is reflected in the low IoU value.
F.4 Efficiency of parameter generation
@@
Appendix G Generalization on Other Tasks
@@
G.1 Object detection
Faster R-CNN [56] utilizes a region proposal network (RPN) which shares full-image convolutional features with the detection network to improve Fast R-CNN on object detection task. The FCOS (Fully Convolutional One-Stage) [69] model is a single-stage object detection model that simplifies the detection process by eliminating the need for anchor boxes. In the object detection task, We implement Faster R-CNN [56] and FCOS [69] with ResNet-50 backbone on the COCO [41] dataset based on torch/torchvision. Considering the time cost of data for p-diff, we directly use the pre-trained parameters as our first training data, then fine-tune it to obtain other training data. The parameters of the boxing predictor layer are generated by p-diff. We report the results in Tab. 19. Our method can get models with similar or even better performance than the original model in seconds.
@@ G.1 Object detection
G.2 Semantic segmentation
Fully Convolutional Network (FCN) [43] was designed to efficiently process and analyze images at the pixel level, allowing for the semantic segmentation of objects within an image. Following the approach in object detection, we implement semantic segmentation task using FCN [43] with ResNet-50 backbone to evaluate a subset of COCO val2017, on the 20 categories that are present in the Pascal VOC dataset. We generate a subset of the parameters of backbone and report the results in Tab. 20. Our approach can generate high-performing neural network parameters in semantic segmentation task.
@@ G.3 Image generation
DDPM [28] is a diffusion-based method in image generation, where UNet [60] is used to model the noise. In the image generation task, we use p-diff to generate a subset of model parameters of UNet. For comparison, we evaluate the p-diff model’s FID score on the CIFAR-10 dataset and report the results in Tab. 21. The best p-diff generated UNet get similar performance to the original model.